assume at the price of 5 naira per unit 50 oranges and 35 pineapple are bought as a result of prices of oranges to 15 naira per unit. the quantity of price of pineapple purchase increase to 70 unit while oranges decreased to 15 unit. what is the cross elasticity for demand of pineapple.
find:
Arc elasticity of pineapple
1
Expert's answer
2019-03-29T10:48:13-0400

The cross elasticity for demand of pineapple is:

Ed = (70 - 35)/(15 - 5)*(15 + 5)/(70 + 35) = 35/10*20/105 = 2/3, so these fruits are substitutes.